A uniform 10 m long, 50 kg pole is lying on the ground. You get your little brother to lift at one end with an upward force of 100N. How far from the other end will you need to lift so the pole stays horizontal (in m)? [Hint: Using your little brother as the fulcrum, what upward force in N, will you need to apply?]

(a) 2.5
(b) 3.8
(c) 5.0
(d) 6.7
(e) 7.5
1
Expert's answer
2020-04-27T10:00:43-0400

The equilibrium conditiions for the pole is

"F_1+F_2=P" and "Pl\/2=F_2(l-x)" then "x=\frac{l(P/2-F_1)}{P-F_1}" hence "x\approx3.8m"

where "l=10m,F_1=100N,P=mg=490N"

(b) 3.8